OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = 5/7x - 12 A. -5x + 7y = -12 B. -5x - 7y = -84 C. -5x + 7y = -84 D. 5x + 7y = -84

OpenStudy (anonymous):

@tanya123

OpenStudy (anonymous):

the answer is C because -5x + 7y = -84, add 5x to both sides, and you get 7y = 5x - 84. so what do you do next @bvbarmy17

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

Hey I'm sorry about that. I had to step out for a bit. Do you still need help with this question?

OpenStudy (anonymous):

well @bvbarmy17 divide by 7

OpenStudy (anonymous):

The equation for standard form is Ax+By=C - Ax cannot be negative - no fractions if possible

OpenStudy (anonymous):

y = 5/7x -12 Subtract Ax from both sides. -5/7x + y = -12 Multiply both sides by 7. 7(-5/7x + y) = 7(-12) = -5x +7y = -84 Ax cannot be negative, so multiply both sides by -1 -1(-5x +7y) = -1(-84) = 5x -7y = 84 But what is odd is that I don't see the correct answer....so I guess you would just back up a step ("the multiplying both sides by -1" step) and go with C. -5x + 7y = -84. But remember Ax should NOT be negative.

OpenStudy (anonymous):

I know my response was late, but I thought I still might try and help.

OpenStudy (anonymous):

lol thanks alot u did help

OpenStudy (anonymous):

Glad to hear that. I shouldn't be away from the computer anymore for at least a few hours, so if there are any more questions, I should be able to help with those.

OpenStudy (anonymous):

awsome thank u alot